### v4

@@ 6.1. Good-for-MDP automata
In general, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. In the syntactic case, the player not only aims to generate a word in $L({\mathcal{N}})$, but also to simultaneously generate an accepting run of ${\mathcal{N}}$ on it. Therefore, any strategy for the syntactic case can be used for the semantic case by ignoring the component of ${\mathcal{N}}$, resulting in a potentially higher semantic satisfaction. Equality, however, is not guaranteed. Indeed, the same way not all automata are history deterministic, it may not be possible to resolve the nondeterminism in ${\mathcal{N}}$ in a way that maximizes the probability of generating a word in $L({\mathcal{N}})$ (thus reaching a maximal semantic satisfaction) while generating rejecting runs on words in $L({\mathcal{N}})$ with zero probability (and thus reaching an equal syntactic satisfaction). When the equality holds for all MDP ${\mathcal{M}}$, we say that ${\mathcal{N}}$ is good-for-MDP (GFM for short). Note that when ${\mathcal{N}}$ is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ×${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability. Note that $\mu^{*}_{\mathcal{N}}$ can be viewed as a pure finite-memory strategy for ${\mathcal{M}}$ for which $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$.
@@ 6.2. The nondeterminism hierarchy and GFMness
Theorem 14.
Every HD automaton is GFM [KMBK14].
@@
Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{N}}$. Also, as $A_{\sigma}=\left\{\ast\right\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{N}}\rightarrow Q_{\mathcal{N}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{N}}(q,\sigma)$ for all $\left\langle\sigma,q\right\rangle\in\Sigma\times Q_{\mathcal{N}}$. Thus, $\mu^{*}$ induces a pruning ${\mathcal{N}}_{\mu^{*}}$ of ${\mathcal{N}}$, and $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{N}}_{\mu^{*}}$.
Now, since $\mu^{*}$ is an optimal strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ and ${\mathcal{N}}$ is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{N}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{N}})$. That is, ${\mathcal{N}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{N}}$ that loses only a set of words of measure zero and hence ${\mathcal{N}}$ is almost-DBP.
Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP. By Theorem 15, we can conclude that the SD-NCW described there, is also not GFM. Thus, while HDness implies GFMness for all accaptence conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness. In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM. In particular, Theorem 15 suggests an alternative proof of Theorem 10, which states that all SD-NBWs are almost-DBP.
@@
Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in ${\mathcal{M}}$ according to $\mu$. Note that since $\mu$ uses $M$ as a memory structure, we may think of ${\mathcal{C}}$ as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise. Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to ${\mathcal{C}}$. Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of ${\mathcal{N}}$ over the word $\tau(r)$ that is sampled when playing in ${\mathcal{M}}$ with the startegy $\mu$. I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
@@
Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of ${\mathcal{M}}$ according to $\mu$ and in addition generates a random run of ${\mathcal{N}}$ over $\tau(r)$. It is not hard to see that ${\mathcal{C}}_{\mathcal{N}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{N}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$. We proved that there is a zero probability that $\tau(r)\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting. Thus, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$.
@@
The co-Büchi case follows from Theorems 11 and 15. For Büchi automata, consider an SD-NBW ${\mathcal{N}}=\left\langle\Sigma,Q_{\mathcal{N}},q_{0},\delta_{\mathcal{N}},\alpha_{\mathcal{N}}\right\rangle$, and a $\Sigma$-labeled MDP ${\mathcal{M}}$. Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ given by [CJH03]. Then by Proposition 16 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$. Since $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ holds for all ${\mathcal{N}}$ and ${\mathcal{M}}$, we have the equality $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
@@
To conclude the GFMness of ${\mathcal{N}}$, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$. Recall that a strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in ${\mathcal{M}}$ and in ${\mathcal{N}}$ simultaneously. In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in ${\mathcal{M}}$, while in the latter it should choose a $\tau(s)$-succsessor of the current state $q$ in ${\mathcal{N}}$. The strategy $\mu^{\prime}$ resolves the first according to to $\mu$, and hence ignores the ${\mathcal{N}}$ component. Then, for the ${\mathcal{N}}$ component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of ${\mathcal{N}}$ on $w_{\mathcal{C}}$ as follows. The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in ${\mathcal{C}}$; in particular, if the ergodic SCC that is reached in ${\mathcal{C}}$ contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$. In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
As the MC ${\mathcal{C}}$ is finite, then with probability 1 an ergodic component of ${\mathcal{C}}$ is reached and all of its states are traversed infinitely often. Hence, with probability 1, we have one of the following outcomes. Either the ergodic component that is reached does not contains an $a$-labeled state, or it contains an $a$-labeled state. In the former case, the run $r_{\mathcal{N}}$ is accepting. Indeed, according to $\mu^{\prime}$, the run $r_{\mathcal{N}}$ moves to $q_{acc}$ once the ergodic component that has no $a$-labeled states is reached, and thus $r_{\mathcal{N}}$ eventually gets stuck at $q_{acc}$ while reading a suffix of $w_{\mathcal{C}}$ that has no $a$’s. In the latter case, $r_{\mathcal{N}}$ is stuck in $q_{0}$ and is thus rejecting, and with probability 1 $w_{\mathcal{C}}$ has infinitely many $a$’s. Hence, in the latter case there is a zero probability that $w_{\mathcal{C}}\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting, and we are done.

### v5

@@ 6.1. Good-for-MDP automata
In general, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. In the syntactic case, the player not only aims to generate a word in $L({\mathcal{N}})$, but also to simultaneously generate an accepting run of $\mathcal{N}$} on it. Therefore, any strategy for the syntactic case can be used for the semantic case by ignoring the component of $\mathcal{N}$}, resulting in a potentially higher semantic satisfaction. Equality, however, is not guaranteed. Indeed, the same way not all automata are history deterministic, it may not be possible to resolve the nondeterminism in $\mathcal{N}$} in a way that maximizes the probability of generating a word in $L({\mathcal{N}})$ (thus reaching a maximal semantic satisfaction) while generating rejecting runs on words in $L({\mathcal{N}})$ with zero probability (and thus reaching an equal syntactic satisfaction). When the equality holds for all MDP $\mathcal{M}$}, we say that $\mathcal{N}$} is good-for-MDP (GFM for short). Note that when $\mathcal{N}$} is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability. Note that $\mu^{*}_{\mathcal{N}}$ can be viewed as a pure finite-memory strategy for $\mathcal{M}$} for which $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$.
@@ 6.2. The nondeterminism hierarchy and GFMness
{thmC}
[[KMBK14]] Every HD automaton is GFM.
@@
Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ${\mathcal{M}}\times{\mathcal{N}}$. Also, as $A_{\sigma}=\left\{\ast\right\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{N}}\rightarrow Q_{\mathcal{N}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{N}}(q,\sigma)$ for all $\left\langle\sigma,q\right\rangle\in\Sigma\times Q_{\mathcal{N}}$. Thus, $\mu^{*}$ induces a pruning ${\mathcal{N}}_{\mu^{*}}$ of $\mathcal{N}$}, and $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{N}}_{\mu^{*}}$.
Now, since $\mu^{*}$ is an optimal strategy for ${\mathcal{M}}\times{\mathcal{N}}$ and $\mathcal{N}$} is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{N}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{N}})$. That is, ${\mathcal{N}}_{\mu^{*}}$ is a deterministic pruning of $\mathcal{N}$} that loses only a set of words of measure zero and hence $\mathcal{N}$} is almost-DBP.
Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP. By Theorem 14, we can conclude that the SD-NCW described there, is also not GFM. Thus, while HDness implies GFMness for all acceptance conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness. In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM. In particular, Theorem 14 suggests an alternative proof of Theorem 10, which states that all SD-NBWs are almost-DBP.
@@
Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in $\mathcal{M}$} according to $\mu$. Note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise. Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to $\mathcal{C}$}. Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of $\mathcal{N}$} over the word $\tau(r)$ that is sampled when playing in $\mathcal{M}$} with the strategy $\mu$. I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
@@
Let $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ be the strategy for ${\mathcal{M}}\times{\mathcal{N}}$ that resolves the actions of $\mathcal{M}$} according to $\mu$ and in addition generates a random run of $\mathcal{N}$} over $\tau(r)$. It is not hard to see that ${\mathcal{C}}_{\mathcal{N}}$ induces the same distribution over infinite paths in $(S\times Q_{\mathcal{N}})^{\omega}$ as the Markov chain ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$. We proved that there is a zero probability that $\tau(r)\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting. Thus, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$.
@@
The co-Büchi case follows from Theorems 11 and 14. For Büchi automata, consider an SD-NBW ${\mathcal{N}}=\left\langle\Sigma,Q_{\mathcal{N}},q_{0},\delta_{\mathcal{N}},\alpha_{\mathcal{N}}\right\rangle$, and a $\Sigma$-labeled MDP $\mathcal{M}$}. Let $\mu$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ given by [CJH03]. Then by Proposition 15 it holds that there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, and we get $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$. Since $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ holds for all $\mathcal{N}$} and $\mathcal{M}$}, we have the equality $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$.
@@
To conclude the GFMness of $\mathcal{N}$}, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$. Recall that a strategy for ${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in $\mathcal{M}$} and in $\mathcal{N}$} simultaneously. In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}. The strategy $\mu^{\prime}$ resolves the first according to $\mu$, and hence ignores the $\mathcal{N}$} component. Then, for the $\mathcal{N}$} component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of $\mathcal{N}$} on $w_{\mathcal{C}}$ as follows. The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in $\mathcal{C}$}; in particular, if the ergodic SCC that is reached in $\mathcal{C}$} contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$. In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
As the MC $\mathcal{C}$} is finite, then with probability 1 an ergodic component of $\mathcal{C}$} is reached and all of its states are traversed infinitely often. Hence, with probability 1, we have one of the following outcomes. Either the ergodic component that is reached does not contain an $a$-labeled state, or it contains an $a$-labeled state. In the former case, the run $r_{\mathcal{N}}$ is accepting. Indeed, according to $\mu^{\prime}$, the run $r_{\mathcal{N}}$ moves to $q_{acc}$ once the ergodic component that has no $a$-labeled states is reached, and thus $r_{\mathcal{N}}$ eventually gets stuck at $q_{acc}$ while reading a suffix of $w_{\mathcal{C}}$ that has no $a$’s. In the latter case, $r_{\mathcal{N}}$ is stuck in $q_{0}$ and is thus rejecting, and with probability 1 $w_{\mathcal{C}}$ has infinitely many $a$’s. Hence, in the latter case there is a zero probability that $w_{\mathcal{C}}\in L({\mathcal{N}})$ and $r_{\mathcal{N}}$ is rejecting, and we are done.